# ExDDV: A New Dataset for Explainable Deepfake Detection in Video

### v1

@@ 4.1 Baseline architectures
BLIP-2. The original BLIP architecture is designed to train a unified VLM end-to-end. It bootstraps from noisy image-text pairs by generating synthetic captions and filtering out noise. The training regime combines contrastive learning, image-text matching and autoregressive language modeling. BLIP-2 takes a different route by avoiding an end-to-end approach. Instead, it introduces a lightweight Querying Transformer (Q-Former) that serves as a bridge between the image encoder and the LLM. The Q-Former employs a fixed set of learnable query tokens to extract the most relevant features from the image encoder, resulting in an efficient design with minimal computational costs. We opt for the BLIP-2 model based on an Open Pretrained Transformer (OPT) backbone with $2.7$B parameters. For efficiency, we use the version optimized through int8 quantization.
Phi-3-Vision. The Phi-3 [1] suite of models ranges in functionality from solving pure language tasks to multimodal (vision and language) tasks. For our task, we employ the Phi-3-Vision model, which adopts a data-centric approach rather than simply increasing model size or applying distinct training regimes. The model is pre-trained using a curated pipeline that combines heavily filtered public web data with synthetic data generated by LLMs. This pre-training strategy first develops a broad language understanding, then enhances the model’s reasoning and logical inference capabilities. We select the architecture based on $128$K tokens, which uses flash attention.
LLaVA-1.5. The LLaVA [45] family seamlessly integrates visual inputs with an LLM to generate natural language descriptions for visual tasks. When an image is provided as input, a pre-trained vision encoder extracts the relevant features, which are then projected into the LLM’s token space through a linear transformation that preserves both spatial and semantic information. The LLM processes this sequence auto-regressively, producing responses that are inherently grounded in visual input. In our application, this approach enables the model not only to accurately answer questions about deepfake detection, but also to provide detailed and interpretable explanations of its reasoning. We opt for the LLaVA-1.5 version, with 7B parameters.
@@ 4.2 In-context learning
Pipeline. In Figure 3, we illustrate our few-shot in-context learning pipeline. The training examples are chosen by a k-nearest neighbor (k-NN) model for each test video. The k-NN model extracts feature vectors with the CLIP [63] image encoder, which is based on the ResNet [30] architecture. Features are only extracted from the training frames that are annotated with clicks. The frames are further used independently by the k-NN. For efficient inference, we store the training feature embeddings along with the associated annotations. When a test image is provided, we obtain its features via a similar process, extracting the ResNet features for the middle frame. Then, we search for the closest $k$ training frames based on the cosine similarity between feature embeddings. The annotations of the nearest neighbors are used to enrich the custom prompt given to the VLM. The k-NN procedure is supposed to provide a set of $k$ deepfake explanations that are potentially relevant for the test video. The custom prompt instructs the VLM to provide a similar explanation for the input test video.
Hyperparameters. We consider two alternative CLIP image encoders for the k-NN: ResNet-50 and ResNet-101. For the number of neighbors $k$, we test values in the set $\{1,3,5,10\}$. The best choice is $k=5$. To perform the k-NN search at various representation levels, we extract features at three different depth levels. We use low-level features right before the first residual block, mid-level features from an intermediate residual block (equally distanced from the input and output), and high-level features from the last layer. The features producing the best results are the high-level ones. We keep the default parameters of all VLMs during inference, except for the temperature, which is set to 00 to reduce the chance of generating hallucinations.
@@ 4.3 Fine-tuning
Pipeline. On the one hand, we follow the default fine-tuning methodology for the BLIP-2 model. On the other hand, given the increased memory requirements of Phi-3-Vision and LLaVA-1.5, we employ Low-Rank Adaptation (LoRA) [31] to fine-tune these two models.
@@
Hyperparameters. Phi-3-Vision and LLaVA are trained for 10 epochs with mini-batches of $16$ and $32$ samples, respectively. BLIP-2 is also fine-tuned for 10 epochs with a mini-batch size of $16$, but with a gradient accumulation of $2$. All models are optimized with AdamW, using a learning rate of $2\cdot 10{-4}$ and a cosine annealing scheduler to reduce the learning rate during training. For Phi-3-V, we use a LoRA rank of $64$ and a dropout rate of $0.05$. For LLaVA, we set the rank to $128$ and do not employ dropout. The parameter $\alpha$ of LoRA is set to $256$ for both models.
@@ 4.4 Click supervision
We harness the additional supervision signal in our dataset, namely the locations of the visual flaws represented by mouse clicks. While this idea was found to be useful in a number of vision tasks [59, 7, 14, 16, 49], to our knowledge, it has not been explored to train explainable video models. The rationale behind using click supervision is to help the explainable model to focus on the region of interest (ROI) and consequently provide precise explanations.
Figure 4: Our click supervision pipeline at inference time. A ViT-based click predictor estimates click coordinates for the input frames. A hard or soft masking is applied to mask the area outside the region of interest. The masked frames are given as input to a VLM. Best viewed in color.
@@
BLIP-2 in-context $0.44_{\pm 0.008}$ $0.19_{\pm 0.004}$ $0.03_{\pm 0.001}$ $0.21_{\pm 0.001}$ $0.14_{\pm 0.003}$ $0.06_{\pm 0.001}$ $0.14_{\pm 0.001}$
BLIP-2 in-context hard $0.47_{\pm 0.002}$ $0.20_{\pm 0.001}$ $0.04_{\pm 0.001}$ $0.21_{\pm 0.003}$ $0.14_{\pm 0.005}$ $0.06_{\pm 0.001}$ $0.14_{\pm 0.002}$
BLIP-2 in-context soft $0.46_{\pm 0.004}$ $0.20_{\pm 0.003}$ $0.04_{\pm 0.001}$ $0.20_{\pm 0.001}$ $0.14_{\pm 0.001}$ $0.06_{\pm 0.001}$ $0.14_{\pm 0.002}$
BLIP-2 fine-tuned $0.45_{\pm 0.009}$ $0.29_{\pm 0.002}$ $0.09_{\pm 0.010}$ $0.14_{\pm 0.015}$ $0.21_{\pm 0.010}$ $0.09_{\pm 0.007}$ $0.20_{\pm 0.009}$
@@
Phi-3-V in-context $0.30_{\pm 0.002}$ $0.14_{\pm 0.003}$ $0.03_{\pm 0.002}$ $0.18_{\pm 0.005}$ $0.15_{\pm 0.002}$ $0.05_{\pm 0.002}$ $0.13_{\pm 0.002}$
Phi-3-V in-context hard $0.30_{\pm 0.002}$ $0.14_{\pm 0.005}$ $0.03_{\pm 0.004}$ $0.18_{\pm 0.005}$ $0.16_{\pm 0.003}$ $0.05_{\pm 0.003}$ $0.13_{\pm 0.005}$
Phi-3-V in-context soft $0.30_{\pm 0.006}$ $0.14_{\pm 0.003}$ $0.03_{\pm 0.004}$ $0.18_{\pm 0.004}$ $0.15_{\pm 0.002}$ $0.05_{\pm 0.003}$ $0.13_{\pm 0.004}$
Phi-3-V fine-tuned $0.42_{\pm 0.005}$ $0.30_{\pm 0.002}$ $0.06_{\pm 0.001}$ $0.20_{\pm 0.004}$ $0.21_{\pm 0.005}$ $0.07_{\pm 0.005}$ $0.19_{\pm 0.006}$
@@
LLaVA in-context $0.48_{\pm 0.002}$ $0.22_{\pm 0.008}$ $0.03_{\pm 0.001}$ $0.21_{\pm 0.001}$ $0.12_{\pm 0.002}$ $0.03_{\pm 0.002}$ $0.10_{\pm 0.002}$
LLaVA in-context hard $0.47_{\pm 0.002}$ $0.29_{\pm 0.010}$ $0.05_{\pm 0.001}$ $0.22_{\pm 0.001}$ $0.20_{\pm 0.002}$ $0.04_{\pm 0.000}$ $0.16_{\pm 0.002}$
LLaVA in-context soft ${\color[rgb]{0,0,1}0.50}_{\pm 0.006}$ $0.23_{\pm 0.004}$ $0.04_{\pm 0.001}$ $0.20_{\pm 0.002}$ $0.12_{\pm 0.002}$ $0.04_{\pm 0.001}$ $0.11_{\pm 0.002}$
LLaVA fine-tuned $0.45_{\pm 0.003}$ $0.33_{\pm 0.003}$ $0.07_{\pm 0.002}$ $0.23_{\pm 0.003}$ $0.24_{\pm 0.001}$ $0.08_{\pm 0.001}$ $0.21_{\pm 0.001}$
@@
In Figure 4, we present our pipeline based on click-supervision, which comprises a click predictor based on ResNet, and a frame masking heuristic, which preserves the ROI and masks the rest of the frame. To train the click predictor, we first extract all the training frames in which a visual glitch is annotated through a click, along with the coordinates of each click. The coordinates are further normalized between 0 and 1 to make the click predictor invariant to distinct frame resolutions. Next, we train a ViT-based regression model to predict click coordinates. The ViT [22] backbone is pre-trained on ImageNet. During inference, we apply the regressor on each test frame to predict the clicks. Then, we mask the image area outside the ROI, which is defined as a round area of radius $r$ around the click coordinates. We consider two alternative masking operations, soft and hard. Hard masking implies replacing the masked pixels with zero. Soft masking is based on a 2D Gaussian distribution centered in the predicted click location, where the $\sigma=r$. The masking operation is performed after the input images are normalized.
5 Experiments and Results
@@
Evaluation measures. Although there are many measures to assess either the semantic similarity or the n-gram overlap of two text samples, such measures are not able to fully capture the similarity between texts. As a result, we evaluate the XAI models using a wide range of metrics, aiming to provide an extensive evaluation of the results. For semantic understanding, we employ Sentence-BERT [64] to embed both predicted and ground-truth descriptions, and compute the cosine similarity between the two. We also employ BERTScore [80] to compute the similarity at the token level. Every token in the ground-truth is greedily matched with a token in the prediction to compute a recall. Similarly, every token in the prediction is matched with a token in the ground-truth to compute the precision. These are combined into an F1 score, called BERTScore. To assess n-gram overlaps, we adopt the most popular evaluation measures used in image captioning: BLEU [60], METEOR [5] and ROUGE [42]. For these metrics, we set the maximum n-gram length to $n=2$. We evaluate click predictors in terms of the mean absolute error (MAE). We run each experiment three times and report the average scores and the standard deviations.
Figure 5: Qualitative examples for LLaVA based on in-context learning and LLaVA fine-tuned with soft input masking.
Quantitative results. The results of our experiments are shown in Table 5. We present the results for all models (BLIP-2, Phi-3-Vision and LLaVA-1.5) and learning scenarios (pre-training, in-context learning and fine-tuning). Click supervision is integrated via soft or hard masking, respectively. Consistent with the inter-annotator agreement scores, we observe that all models yield better scores in terms of semantic measures than n-gram overlap measures. The gap can be explained by the fact that models generate varied outputs, often using alternative phrases and words to express the same concept.
@@
The reported scores also attest the advantages of integrating click predictions via soft and hard masking. This is observable for all three models, although the gains are somewhat lower for LLaVA. Both masking strategies appear to be equally effective. In response to RQ2, we find that the benefits of using click supervision has a positive impact, boosting the performance of all tested VLMs.
In Figure 6, we showcase the performance of the fine-tuned LLaVA model for different dimensions of the training set. The performance reaches a plateau after 2,000 training samples. This observation suggests that our dataset contains enough samples to train an explainable deepfake detection model, ellucidating RQ3.
Figure 6: The performance of fine-tuned LLaVA against the number of samples used for training.
In Table 6, we report the mean absolute errors for both ViT and ResNet-50 click predictors. The error of ViT is slightly lower, representing an average offset of only $12$ pixels w.r.t. the ground-truth coordinates. These results confirm that the regression models can accurately localize visual artifacts, thus providing a positive answer to RQ4.
Qualitative results. Besides the quantitative measurements, we also present qualitative results. In Figure 5, we illustrate some explanations provided by various variants of LLaVA. The examples include both relevant explanations as well as wrong explanations, e.g. identifying artifacts on real videos. We present examples for additional versions of LLaVA in Figure 8 from the supplementary.
@@
In Figure 10 from the supplementary, we showcase some examples of how the ViT-based click predictor compares with the ground-truth click locations. We observe that the predictor is able to precisely locate visual artifacts.

### v2

@@ 4.1 Baseline architectures
BLIP-2. The original BLIP architecture is designed to train a unified VLM end-to-end. It bootstraps from noisy image-text pairs by generating synthetic captions and filtering out noise. The training regime combines contrastive learning, image-text matching and autoregressive language modeling. BLIP-2 avoids end-to-end learning by taking a different route. More specifically, it introduces a lightweight Querying Transformer (Q-Former) that serves as a bridge between the image encoder and the LLM. The Q-Former employs a fixed set of learnable query tokens to extract the most relevant features from the image encoder, resulting in an efficient design with minimal computational costs. We opt for the BLIP-2 model based on an Open Pretrained Transformer (OPT) backbone with 2.7B parameters. For efficiency, we use the version optimized through int8 quantization.
Phi-3-Vision. The Phi-3 [abdin-arxiv-2024] suite of models ranges in functionality from solving pure language tasks to multimodal (vision and language) tasks. For our task, we employ the Phi-3-Vision model, which adopts a data-centric approach rather than simply increasing model size or applying distinct training regimes. The model is pre-trained using a curated pipeline that combines heavily filtered public web data with synthetic data generated by LLMs. This pre-training strategy first develops a broad language understanding, then enhances the model’s reasoning and logical inference capabilities. We select the architecture based on 128K tokens, which uses flash attention.
LLaVA-1.5. The LLaVA [liu-NeurIPS-2023] family seamlessly integrates visual inputs with an LLM to generate natural language descriptions for visual tasks. When an image is provided as input, a pre-trained vision encoder extracts the relevant features. These are further projected into the LLM’s token space through a linear transformation that preserves both spatial and semantic information. The LLM processes this sequence auto-regressively, producing responses that are inherently grounded in visual input. In our application, this approach enables the model not only to accurately answer questions about deepfake detection, but also to provide detailed and interpretable explanations of its reasoning. We opt for the LLaVA-1.5 version, with 7B parameters.
@@ 4.2 In-context learning
Pipeline. In Fig. 3, we illustrate our few-shot in-context learning pipeline. The training examples are chosen by a k-nearest neighbor (k-NN) model for each test video. The k-NN model extracts feature vectors with the CLIP [radford-PmLR-2021] image encoder, which is based on the ResNet [he-CVPR-2016] architecture. Features are only extracted from the training frames that are annotated with clicks. The frames are further used independently by the k-NN. For efficient inference, we store the training feature embeddings along with the associated annotations. When a test image is provided, we obtain its features via a similar process, extracting the ResNet features for the middle frame. Then, we search for the closest $k$ training frames based on the cosine similarity between feature embeddings. The annotations of the nearest neighbors are used to enrich the custom prompt given to the VLM. The k-NN procedure is supposed to provide a set of $k$ deepfake explanations that are potentially relevant for the test video. The custom prompt instructs the VLM to provide a similar explanation for the input test video.
Hyperparameters. We consider two alternative CLIP image encoders for the k-NN: ResNet-50 and ResNet-101. For the number of neighbors $k$, we test values in the set $\{1,3,5,10\}$. The best choice is $k=5$. To perform the k-NN search at various representation levels, we extract features at three different depth levels. We use low-level features right before the first residual block, mid-level features from an intermediate residual block (equally distanced from the input and output), and high-level features from the last layer. The features producing the best results are the high-level ones. We keep the default parameters of all VLMs during inference, except for the temperature, which is set to 0 to reduce the chance of generating hallucinations.
@@ 4.3 Fine-tuning
Pipeline. On the one hand, we follow the default fine-tuning methodology for the BLIP-2 model. On the other hand, given the increased memory requirements of Phi-3-Vision and LLaVA-1.5, we employ Low-Rank Adaptation (LoRA) [hu-ICLR-2022] to fine-tune these two models.
@@
Figure 4: Our click supervision pipeline at inference time. A ViT-based click predictor estimates click coordinates for the input frames. A hard or soft masking is applied to mask the area outside the region of interest. The masked frames are given as input to a fine-tuned VLM. Best viewed in color.
Hyperparameters. Phi-3-Vision and LLaVA are trained for 10 epochs with mini-batches of $16$ and $32$ samples, respectively. BLIP-2 is also fine-tuned for 10 epochs with a mini-batch size of $16$, but with a gradient accumulation of $2$. All models are optimized with AdamW, using a learning rate of $2\cdot 10^{-4}$ and a cosine annealing scheduler to reduce the learning rate during training. For Phi-3-V, we use a LoRA rank of $64$ and a dropout rate of $0.05$. For LLaVA, we set the rank to $128$ and do not employ dropout. The parameter $\alpha$ of LoRA is set to $256$ for both models.
@@ 4.4 Click supervision
We harness the additional supervision signal in our dataset, namely the locations of the visual flaws represented by mouse clicks. While this idea was found to be useful in a number of vision tasks [papadopoulos-CVPR-2017, benenson-CVPR-2019, chen-ECCV-2022, cheng-CVPR-2022, luo-CVPR-2024], to our knowledge, it has not been explored to train explainable video models. The rationale behind using click supervision is to guide the explainable model towards focusing on the region of interest (ROI) and consequently provide precise explanations.
In Fig. 4, we present our pipeline based on click-supervision, which comprises a click predictor based on ResNet, and a frame masking heuristic, which preserves the ROI, while masking the rest of the frame. To train the click predictor, we first extract all the training frames in which a visual glitch is annotated through a click, along with the coordinates of each click. The coordinates are further normalized between 0 and 1 to make the click predictor invariant to distinct frame resolutions. Next, we train a ViT-based regression model to predict click coordinates. The ViT [dosovitskiy-ICLR-2021] backbone is pre-trained on ImageNet. During inference, we apply the regressor on each test frame to predict the clicks. Then, we mask the image area outside the ROI, which is defined as a round area of radius $r$ around the click coordinates. We consider two alternative masking operations, soft and hard. Hard masking implies replacing the masked pixels with zero. Soft masking is based on a 2D Gaussian distribution centered in the predicted click location, where the $\sigma=r$. The masking operation is performed after the input images are normalized.
5 Experiments and Results
@@
BLIP-2 in-context $0.44_{\pm 0.008}$ $19_{\pm 0.004} 0.0$$03_{\pm 0.001}$ $0.21_{\pm 0.001}$ $14_{\pm 0.003} 0.0$$06_{\pm 0.001}$ $0.14_{\pm 0.001}$
BLIP-2 in-context hard $0.47_{\pm 0.002}$ $20_{\pm 0.001} 0.0$$04_{\pm 0.001}$ $0.21_{\pm 0.003}$ $14_{\pm 0.005} 0.0$$06_{\pm 0.001}$ $0.14_{\pm 0.002}$
BLIP-2 in-context soft $0.46_{\pm 0.004}$ $20_{\pm 0.003} 0.0$$04_{\pm 0.001}$ $0.20_{\pm 0.001}$ $14_{\pm 0.001} 0.0$$06_{\pm 0.001}$ $0.14_{\pm 0.002}$
BLIP-2 fine-tuned $0.45_{\pm 0.009}$ $29_{\pm 0.002} 0.0$$09_{\pm 0.010}$ $0.14_{\pm 0.015}$ $21_{\pm 0.010} 0.0$$09_{\pm 0.007}$ $0.20_{\pm 0.009}$
@@
Phi-3-V in-context $0.30_{\pm 0.002}$ $14_{\pm 0.003} 0.0$$03_{\pm 0.002}$ $0.18_{\pm 0.005}$ $15_{\pm 0.002} 0.0$$05_{\pm 0.002}$ $0.13_{\pm 0.002}$
Phi-3-V in-context hard $0.30_{\pm 0.002}$ $14_{\pm 0.005} 0.0$$03_{\pm 0.004}$ $0.18_{\pm 0.005}$ $16_{\pm 0.003} 0.0$$05_{\pm 0.003}$ $0.13_{\pm 0.005}$
Phi-3-V in-context soft $0.30_{\pm 0.006}$ $14_{\pm 0.003} 0.0$$03_{\pm 0.004}$ $0.18_{\pm 0.004}$ $15_{\pm 0.002} 0.0$$05_{\pm 0.003}$ $0.13_{\pm 0.004}$
Phi-3-V fine-tuned $0.42_{\pm 0.005}$ $30_{\pm 0.002} 0.0$$06_{\pm 0.001}$ $0.20_{\pm 0.004}$ $21_{\pm 0.005} 0.0$$07_{\pm 0.005}$ $0.19_{\pm 0.006}$
@@
LLaVA in-context $0.48_{\pm 0.002}$ $22_{\pm 0.008} 0.0$$03_{\pm 0.001}$ $0.21_{\pm 0.001}$ $12_{\pm 0.002} 0.0$$03_{\pm 0.002}$ $0.10_{\pm 0.002}$
LLaVA in-context hard $0.47_{\pm 0.002}$ $29_{\pm 0.010} 0.0$$05_{\pm 0.001}$ $0.22_{\pm 0.001}$ $20_{\pm 0.002} 0.0$$04_{\pm 0.000}$ $0.16_{\pm 0.002}$
LLaVA in-context soft ${\color[rgb]{0,0,1}0.50}_{\pm 0.006}$ $23_{\pm 0.004} 0.0$$04_{\pm 0.001}$ $0.20_{\pm 0.002}$ $12_{\pm 0.002} 0.0$$04_{\pm 0.001}$ $0.11_{\pm 0.002}$
LLaVA fine-tuned $0.45_{\pm 0.003}$ $33_{\pm 0.003} 0.0$$07_{\pm 0.002}$ $0.23_{\pm 0.003}$ $24_{\pm 0.001} 0.0$$08_{\pm 0.001}$ $0.21_{\pm 0.001}$
@@
Evaluation measures. Although there are many measures to assess either the semantic similarity or the n-gram overlap of two text samples, such measures are not able to fully capture the similarity between texts. As a result, we evaluate the XAI models using a wide range of metrics, aiming to provide an extensive evaluation of the results. For semantic understanding, we employ Sentence-BERT [reimers-EMNLP-2019] to embed both predicted and ground-truth descriptions, and compute the cosine similarity between the two. We also employ BERTScore [zhang-ICLR-2020] to compute the similarity at the token level. Every token in the ground-truth is greedily matched with a token in the prediction to compute a recall. Similarly, every token in the prediction is matched with a token in the ground-truth to compute the precision. These are combined into an F1 score, called BERTScore. To assess n-gram overlaps, we adopt the most popular evaluation measures used in image captioning: BLEU [Papineni-ACL-2002], METEOR [banerjee-ACL-2005] and ROUGE [lin-ACL-2004]. For these metrics, we set the maximum n-gram length to $n=2$. We evaluate click predictors in terms of the mean absolute error (MAE). We run each experiment three times and report the average scores and the standard deviations.
Quantitative results. The results of our experiments are shown in Tab. 5. We present the results for all models (BLIP-2, Phi-3-Vision and LLaVA-1.5) and learning scenarios (pre-training, in-context learning and fine-tuning). Click supervision is integrated via soft or hard masking, respectively. Consistent with the inter-annotator agreement scores, we observe that all models yield better scores in terms of semantic measures than n-gram overlap measures. The gap can be explained by the fact that models generate varied outputs, often using alternative phrases and words to express the same concept.
Figure 5: The performance of fine-tuned LLaVA (vertical axis) against the number of samples used for training (horizontal axis).
Figure 6: Qualitative examples for LLaVA based on in-context learning and LLaVA fine-tuned with soft input masking. Best viewed in color.
@@
The reported scores also attest the advantages of integrating click predictions via soft and hard masking. This is observable for all three models, although the gains are somewhat lower for LLaVA. Both masking strategies appear to be equally effective. In response to RQ2, we find that using click supervision has a positive impact, boosting the performance of all tested VLMs.
In Fig. 5, we showcase the performance of the fine-tuned LLaVA model for different dimensions of the training set. The performance reaches a plateau after 2,000 training samples. This observation suggests that our dataset contains enough samples to train an explainable deepfake detection model, elucidating RQ3.
In Tab. 6, we report the mean absolute errors for both ViT and ResNet-50 click predictors. The error of ViT is slightly lower, representing an average offset of only $12$ pixels w.r.t. the ground-truth coordinates. These results confirm that the regression models can accurately localize visual artifacts, thus providing a positive answer to RQ4.
@@
Qualitative results. Besides the quantitative measurements, we also present qualitative results. In Fig. 6, we illustrate some explanations provided by two variants of LLaVA. The examples include both relevant explanations as well as wrong explanations, e.g. identifying artifacts on real videos. We present examples for additional versions of LLaVA in Fig. 10 from the supplementary.
In Fig. 7, we showcase some examples of how the ViT-based click predictor compares with the ground-truth click locations. We observe that the predictor is able to precisely locate visual artifacts. We present additional click prediction examples in Fig. 11 from the supplementary.
Figure 7: Qualitative examples of our click predictor. The green dots represent the ground-truth locations and the blue dots represent the corresponding predictions. Best viewed in color.